Sue has $2.30 in dimes and nickels. If she has 11 more dimes than nickels, how many of each coin does she have?
? nickels
? dimes

Answer by mananth(16949)

nickels x
dimes x+11
Value 2.3
5x+10(x +11)= 230
5x+10x+110= 230

15x+110 =230
15x=230 -110
15x=120
/15
x=8 nickels
dimes= 19